# A flexible framework for structural plasticity in GPU-accelerated sparse spiking neural networks

### v1

@@ 1 Introduction
A simulation of a Spiking Neural Network (SNN) is usually a two-step process: First, data structures for neurons and synapses are instantiated on the machine to define the network, including its connectivity; afterwards, the network state is calculated over time starting from some initial conditions according to the neuron and synapse dynamics. This approach assumes that the connectivity is fixed once the network is constructed (typically using population-level connection rules [1]) and at most the strength of connections may be subject to change during the simulation [2]. However, biological evidence suggests that the connectivity itself is continuously changing, not only during development but also in the adult brain [3]. These changes are associated with learning, adaptation, memory, recovery from lesions, drug addiction, and occur following enriched experience, after sensory deprivation or upon stimulation. Besides the biological evidence, it has been found that algorithms abstracted from structural plasticity can benefit machine-learning (ML) models – in particular in the form of pruning [4].
@@
To shed light on candidate mechanisms of structural plasticity and their function for learning and memory, computational models have been developed to study, for example, whether changes are induced by activity, homeostasis, stochasticity, or other factors and simulator developers have implemented functionality for modifying networks during runtime (i.e., forming or eliminating synapses) to simulate these models, e.g. [5, 6, 7, 8, 9, 10, 11]. However, structural plasticity remains underexplored because simulating networks with it is computationally expensive: First, structural changes in biology are slow (hours, days) and therefore need long simulation durations. Second, there are complex inter-dependencies with other biological mechanisms such as synaptic plasticity and some processes may need global instead of only local information, hence limiting implementations. Third, continuously applying changes to the connectivity clashes with the traditional approach of keeping the network structure untouched in computer memory during state propagation which requires to rethink how data structures are handled. Different simulation platforms including neuromorphic systems therefore face different difficulties.
@@
Structural plasticity algorithms have previously largely been implemented in CPU-based SNN simulators [[, e.g.,]]DiazPier2016,Vitay2015 and neuromorphic systems such as SpiNNaker, where the main computational elements are scalar CPU cores [[, e.g.,]]Bogdan2018,Liu2018 and thus the standard routines used for modifying connectivity can be employed. [11] presented a structural plasticity implementation for the BrainScaleS–2 [13] Plasticity Processing Unit (PPU) – a Single Instruction Multiple Thread (SIMD) processor that is tightly integrated with the BrainScaleS–2 neuromorphic system. Here, we propose what we believe is the first structural plasticity framework for implementation on general purpose parallel hardware such as GPUs. The framework aims to combine flexibility (i.e., the possibility to freely explore different structural plasticity mechanisms) with efficiency (i.e., fast simulations). We identify the following constraints:
@@
The GeNN simulator [14, 15] was originally developed for Computational Neuroscience research but its longstanding focus on flexibility and its efficient implementation of sparse data structures on GPU accelerators has enabled it to become a powerful tool for many areas of SNN research. We have previously added support for various machine-learning-specific functionalities including parallel batch simulation of models – allowing multiple copies of the model to be run simultaneously in order to maximize GPU occupancy – and user-defined ‘custom update’ operations [16, 17]. Here, we present a new extension to GeNN which allows users to easily define hybrid CPU/GPU structural plasticity algorithms using the same code-generation based approach already used to define neuron models and learning rules as well as initialization algorithms.
Using this framework we demonstrate two applications. Firstly, building on our previous work [17, 18] using the e-prop learning rule [19] to efficiently train recurrent SNN classifiers, we combine the GeNN implementation of e-prop with the DEEP R structural plasticity rule [20] – implemented in our new framework – to train very sparse networks on the DVS gesture classification task [21] with no loss of accuracy while using $90\times$ fewer parameters than comparable recurrent dense SNN models. The time overhead of applying structural plasticity is less than $1\,\mathrm{\%}$ and the sparse models used are an order of magnitude faster to train than their dense counterparts. Secondly, we implement a model of topographic map formation [22] in GeNN, motivated by a recent example on the SpiNNaker platform [7]. The spatially embedded network employs an activity-independent formation rule: new connections are drawn according to a Gaussian dependence on the lateral distance between neurons. However, synapse elimination is activity-dependent: Spike-Timing-Dependent Plasticity (STDP) [2] varies synaptic strengths and weak synapses are more likely to be removed. The interplay of different plasticity mechanisms in the model results in a refinement of receptive fields and an embedding of learnt input structures in the network connectivity. Even though the large part of the simulation time is here spent on connectivity updates, the simulation is still faster than previous implementations [7] and enables network upscaling experiments to test the scaling performance.
@@ 2.1 GeNN custom connectivity updates
Figure 1: Illustration of data structures used to implement sparse connectivity in GeNN. Five presynaptic neurons are connected to four postsynaptic neurons using sparse connectivity with a maximum row length of three. This is represented in memory as a ragged matrix, where each row n contains the indices of the target neurons of pre-synaptic neuron n. Spikes are propagated through ragged matrix connectivity by parallelizing across the columns as indicated by the snaking lines.
GeNN [14, 15] is a library for generating CUDA [23] code for the simulation of SNN models. GeNN handles much of the complexity of using CUDA and automatically performs device-specific optimizations so as to maximize performance. GeNN consists of a main library – implementing the API used to define models as well as the generic parts of the code generator – and an additional library for each backend (currently there is a reference C++ backend for generating CPU code and a CUDA backend, with an AMD HIP [24] backend under development). PyGeNN [15] exposes these libraries to Python using PyBind11 [25] and provides a Pythonic and user-friendly interface to the low-level functionality provided by GeNN.
One key advantage of GeNN over SNN simulation libraries such as SNNtorch [26], NORSE [27] or JAXsnn [28] – which are built on standard Machine Learning toolkits like PyTorch or JAX – is that GeNN not only supports standard dense (i.e., all-to-all) weight matrices stored as 2D arrays, but also sparse connectivity stored in the ‘ragged matrix’ data structure illustrated in figure 1. In this data structure, each row contains the indices of a single presynaptic neuron’s postsynaptic targets and, to simplify address calculation, rows are padded to the largest number of postsynaptic targets present in the connectivity matrix. Additional 2D arrays with the same shape are used to store synaptic variables such as weights and delays. To efficiently parallelize the propagation of spikes through this data structure, one CUDA thread is allocated to each column of the matrix. Each thread then loops through incoming spikes, with each ‘warps’ of $32$ adjacent threads performing efficient coalesced reads of synaptic weights and accumulating input to the postsynaptic neurons using atomic operations (which on modern GPUs are handled efficiently by the L2 cache). If connectivity updates can be parallelized across presynaptic neurons, this format actually fits the constraints on parallel connectivity updates we defined in the introduction rather well. By adding additional padding to the end of each row, connections can be added and removed without re-allocating the data structure. Because the order of synapses in a row is unimportant, when adding synapses they can simply be added to the end of the row and, when removing them, the synapse at the end of the row can just be moved to the location of the removed synapse and the row-length reduced by one.
@@
DEEP R [20] needs to detect changes to the sign of weights and the synaptic rewiring rule developed by [22] compares weights to a threshold.
@@
DEEP R [20] needs to track the number of synapses which are removed from each row of the synaptic matrix so they can be summed and the same number of new synapses added.
@@
The rule developed by [29] needs to access pre- and postsynaptic calcium variables (essentially low-pass filtered spike trains).
@@
DEEP R [20] needs to be able to sample from a random number generator to create new synapses and the rule developed by [22] uses probabilistic thresholds to determine whether to actually eliminate or add candidate connections.
While, for convenience, GeNN provides some built-in models, one of PyGeNN’s most powerful features is that it enables users to easily define their own models for neurons, synapses and custom updates (which are typically used to implement optimizers and reductions for machine learning models [18]) from within Python using strings containing GeNNCode which is a subset of C99. Continuing this flexible approach, we have added support for a new model primitive – ‘custom connectivity updates’ – to GeNN which provides a framework to build structural plasticity rules. This framework is relatively complex so, while in the remainder of this section we will try to give a flavor of what is possible, we refer interested readers to the GeNN documentation (https://genn-team.github.io) for more details. Rules can be defined with the pygenn.create_custom_connectivity_update_model helper function and can access neuron and synapse variables via ‘variable references’ (pre_var_refs and post_var_refs; and var_refs respectively). Additionally, they provide one GeNNCode string which is run on the GPU and parallelized across the rows of the synaptic matrix (row_update_code) and another that is evaluated serially on the CPU (host_update_code). Within all of these code strings, a suitable random number generator is available via a family of gennrand functions. Within the row_update_code, the index of the presynaptic neuron associated with the row of connectivity can be accessed via the internal variable id_pre and synapses can be added using the add_synapse function which takes the index the postsynaptic target and the initial values of any synaptic variables, in the order they are specified in var_refs. Variables which are associated with the connectivity being updated but are not referenced in this way will be initialized to zero for the new synapse. Using this functionality, we can define a simple example which adds diagonal connections to the connectivity matrix with a synaptic weight (referenced via g) of $1.0$:
@@
At runtime, the update can be triggered (along with any other updates in the “update_connectivity” group) by simply calling model.custom_update("update_connectivity"). If row updates need to remove or otherwise access existing synapses they must iterate through them as rows of the ragged matrix data structure are unsorted. This is implemented by the for_each_synapse statement, within the body of which, the index of current postsynaptic targets can be accessed via the id_post variable (as well as the values of any variables or variable references associated with individual synapses or postsynaptic neurons). Synapses can also be removed within the body of the for_each_synapse statement using the remove_synapse function. Finally, custom connectivity updates can define their own additional neuron and synapse variables (pre_vars and post_vars; and vars respectively). Using this additional functionality, we can define another example which, on the CPU, first loops through each presynaptic neuron and picks a random postsynaptic target to try and remove. Then, on the GPU, we search each row of connectivity in parallel and, if the chosen target is found, remove it:
@@
In order to facilitate postsynaptic spike-triggered updates to synapses, such as those involved in Spike Timing Dependent Plasticity (STDP), GeNN builds an additional transpose matrix data structure (see [30] for a complete description). This data structure is normally built once at initialization time but, if custom connectivity updates are applied to connections with these data structures, they get rebuilt after each custom connectivity update.
@@
While we have previously implemented complex machine learning workflows with networks of this sort in PyGeNN [17, 16], the syntax is designed for computational neuroscience and it can become tedious and difficult to maintain. Therefore, we have built mlGeNN on top of GeNN to provide a Keras-like abstraction [31, 18] with support for both converting pre-trained ANNs to SNNs and for directly training SNNs using e-prop and Eventprop [32]. Here we have further extended mlGeNN to support DEEP R which can be invoked by creating a sparsely-connected network and ‘compiling’ it using an EPropCompiler instantiated as follows:
@@ 2.1.2 DEEP R
Figure 2: Illustration of data structures used for the DEEP R implementation. Five presynaptic neurons are connected to four postsynaptic neurons using sparse connectivity with a maximum row length of three. Each panel shows a network sketch, the ragged matrix and two bitfields – one mirroring the connectivity to provide a faster lookup and the other holding the fixed sign of each potential connection. (A) Initial state. Weights whose sign no longer matches the sign bit are highlighted in green. (B) Elimination pass. Connections highlighted in (A) are removed, the number of connections in each row made ‘dormant’ are counted and the bitfield is updated. (C) Formation pass. Yellow highlighting indicates new the randomly added connections.
The rewiring algorithm DEEP R [20] is designed for training efficient sparse neural networks by replacing weak connections with randomly sampled new ones. DEEP R operates alongside a gradient-based learning rule and marks synapses as ‘dormant’ if weight updates would have led to weights changing sign. Dormant connections are removed and an equal number of randomly chosen synapses are added as replacements. Unlike typical incremental pruning approaches, DEEP R maintains a constant level of connection sparsity throughout training. However, like many implementations of pruning in standard ML tools, the original implementation represents the sparse connectivity as an additional dense weight matrix containing the correct sign of each connection. Here we implement a more efficient version using the structural connectivity framework described in section 2.1 which only requires two additional bitfield data structures on top of the standard GeNN sparse matrix format – one to track the sign and one the presence of each potential synaptic connection.
After the standard initialization of synaptic connectivity, a custom connectivity update is used to initialize the two additional bitfields as illustrated in figure 2A. First, all connectivity bits are cleared and all sign bits are randomized to fix the signs for all potential connections. Then, the update loops through all the initialized synapses in each row, sets the corresponding connectivity bit and ensures the corresponding sign bit matches the sign of the initial weight of the actual connection.
After each training batch, DEEP R is implemented by launching a sequence of custom updates.
@@
is implemented by a final custom connectivity update which first uses the CPU to sum the row-wise dormant synapse count and then re-distributes these uniformly as new ‘activations’ between the rows of synaptic connectivity. Then, on the GPU, this new synapses are created for each activation, checking the connectivity bitfield to ensure there are not duplicate synapses (figure 2B–C).
In its original form [20], DEEP R also features an additional noise term but, following [19] – who first combined DEEP R with e-prop – we do not include this in our implementation.

### v2

@@ 1 Introduction
A simulation of a Spiking Neural Network (SNN) is usually a two-step process: First, data structures for neurons and synapses are instantiated on the machine to define the network, including its connectivity; afterwards, the network state is calculated over time starting from some initial conditions according to the neuron and synapse dynamics. This approach assumes that both neurons and their connectivity are fixed once the network is constructed (typically using population-level connection rules [undef]) and at most the strength of connections may be subject to change during the simulation [undefa]. However, biological evidence suggests that the connectivity itself is continuously changing, not only during development but also in the adult brain [undefb]. These changes are associated with learning, adaptation, memory, recovery from lesions, drug addiction, and occur following enriched experience, after sensory deprivation or upon stimulation. The generation of new neurons in adult brains (neurogenesis) is less ubiquitous, but has also been observed experimentally – in particular in the rodent hippocampal dentate gyrus [undefc, undefd]. Besides the biological evidence, it has been found that algorithms abstracted from structural plasticity can benefit machine-learning (ML) models. Pruning [undefe] reduces the parameter count of large models while preserving accuracy and, inspired by neurogenesis, constructive neural networks [undeff, undefg] are ‘grown’ as part of the training process. . The type of structural plasticity addressed in this work does not add or remove neurons (as in most brain regions beyond development) and concentrates on dynamic formation and elimination of synapses.
@@
To shed light on candidate structural plasticity mechanisms and their function for learning and memory, computational models have been developed to study, for example, whether changes are induced by activity, homeostasis, stochasticity, or other factors. Simulator developers have implemented functionality for modifying network connectivity during runtime to simulate these models, e.g. [undefh, undefi, undefj, undefk, undefl, undefm, undefn]. However, structural plasticity remains underexplored because simulating networks with it is computationally expensive: First, structural changes in biology are slow (hours, days) and therefore need long simulation durations. Second, there are complex inter-dependencies with other biological mechanisms such as synaptic plasticity and some processes may need global instead of only local information, hence limiting implementations. Third, continuously applying changes to the connectivity clashes with the traditional approach of keeping the network structure untouched in computer memory during state propagation which requires to rethink how data structures are handled. Different simulation platforms including neuromorphic systems therefore face different difficulties.
@@
Structural plasticity algorithms have previously largely been implemented in CPU-based SNN simulators [[, e.g.,]]DiazPier2016,Vitay2015 and neuromorphic systems such as SpiNNaker [undefp], where the main computational elements are scalar CPU cores [[, e.g.,]]Bogdan2018,Liu2018 and thus the standard routines used for modifying connectivity can be employed. [undefn] presented a structural plasticity implementation for the BrainScaleS–2 [undefq] Plasticity Processing Unit (PPU) – a Single Instruction Multiple Thread (SIMD) processor that is tightly integrated with the BrainScaleS–2 neuromorphic system. Here, we propose what we believe is the first structural plasticity framework for implementation on general purpose parallel hardware such as GPUs. The framework aims to combine flexibility (i.e., the possibility to freely explore different structural plasticity mechanisms) with efficiency (i.e., fast simulations). We identify the following constraints:
@@
The GeNN simulator [undefr, undefs] was originally developed for Computational Neuroscience research but its longstanding focus on flexibility and its efficient implementation of sparse data structures on GPU accelerators has enabled it to become a powerful tool for many areas of SNN research. We have previously added support for various machine-learning-specific functionalities including parallel batch simulation of models – allowing multiple copies of the model to be run simultaneously in order to maximize GPU occupancy – and user-defined ‘custom update’ operations [undeft, undefu]. Here, we present a new extension to GeNN which allows users to easily define hybrid CPU/GPU structural plasticity algorithms using the same code-generation based approach already used to define neuron models and learning rules as well as initialization algorithms.
Using this framework we demonstrate two applications. Firstly, building on our previous work [undefu, undefv] using the e-prop learning rule [undefw] to efficiently train recurrent SNN classifiers, we combine the GeNN implementation of e-prop with the DEEP R structural plasticity rule [undefx] – implemented in our new framework – to train very sparse networks on the N-MNIST [undefy] and DVS gesture classification tasks [undefz] with no loss of accuracy while using $90\times$ fewer parameters than comparable recurrent dense SNN models. The time overhead of applying structural plasticity is less than $1\,\mathrm{\%}$ and the sparse models used are an order of magnitude faster to train than their dense counterparts. Secondly, we implement a model of topographic map formation [undefaa] in GeNN, motivated by a recent example on the SpiNNaker platform [undefj]. The spatially embedded network employs an activity-independent formation rule: new connections are drawn according to a Gaussian dependence on the lateral distance between neurons. However, synapse elimination is activity-dependent: Spike-Timing-Dependent Plasticity (STDP) [undefa] varies synaptic strengths and weak synapses are more likely to be removed. The interplay of different plasticity mechanisms in the model results in a refinement of receptive fields and an embedding of learnt input structures in the network connectivity. Even though the large part of the simulation time is here spent on connectivity updates, the simulation is still faster than previous implementations [undefj] and enables network upscaling experiments to test the scaling performance.
@@ 2.1 GeNN custom connectivity updates
Figure 1: Illustration of data structures used to implement sparse connectivity in GeNN. Five presynaptic neurons are connected to four postsynaptic neurons using sparse connectivity with a maximum out-degree of two. This is represented in memory as a ragged matrix, where each row n contains the indices of the target neurons of pre-synaptic neuron n. Spikes are propagated through ragged matrix connectivity by parallelizing across the columns as indicated by the snaking lines.
GeNN [undefr, undefs] is a library for generating CUDA [undefab] code for the simulation of SNN models. GeNN handles much of the complexity of using CUDA and automatically performs device-specific optimizations so as to maximize performance. GeNN consists of a main library – implementing the API used to define models as well as the generic parts of the code generator – and an additional library for each backend (currently there is a reference C++ backend for generating CPU code and a CUDA backend, with an AMD HIP [undefac] backend under development). PyGeNN [undefs] exposes these libraries to Python using PyBind11 [undefad] and provides a Pythonic and user-friendly interface to the low-level functionality provided by GeNN.
One key advantage of GeNN over SNN simulation libraries such as SNNtorch [undefae], NORSE [undefaf] or JAXsnn [undefag] – which are built on standard Machine Learning toolkits like PyTorch or JAX – is that GeNN not only supports standard dense (i.e., all-to-all) weight matrices stored as 2D arrays, but also sparse connectivity stored in the ‘ragged matrix’ data structure illustrated in figure 1. In this data structure, each row contains the indices of a single presynaptic neuron’s postsynaptic targets and, to simplify address calculation, rows are padded to the largest number of postsynaptic targets present in the connectivity matrix. Additional 2D arrays with the same shape are used to store synaptic variables such as weights and delays. To efficiently parallelize the propagation of spikes through this data structure, one CUDA thread is allocated to each column of the matrix. Each thread then loops through incoming spikes, with each ‘warp’ of $32$ adjacent threads performing efficient coalesced reads of synaptic weights and accumulating input to the postsynaptic neurons using atomic operations (which on modern GPUs are handled efficiently by the L2 cache). If connectivity updates can be parallelized across presynaptic neurons, this format actually fits the constraints on parallel connectivity updates we defined in the introduction rather well. By adding additional padding to the end of each row, connections can be added and removed without re-allocating the data structure. Because the order of synapses in a row is unimportant, when adding synapses they can simply be added to the end of the row and, when removing them, the synapse at the end of the row can just be moved to the location of the removed synapse and the out-degree reduced by one.
@@
DEEP R [undefx] needs to detect changes to the sign of weights and the synaptic rewiring rule developed by [undefaa] compares weights to a threshold.
@@
DEEP R [undefx] needs to track the number of synapses which are removed from each row of the synaptic matrix so they can be summed and the same number of new synapses added.
@@
The rule developed by [undefah] needs to access pre- and postsynaptic calcium variables (essentially low-pass filtered spike trains).
@@
DEEP R [undefx] needs to be able to sample from a random number generator to create new synapses and the rule developed by [undefaa] uses probabilistic thresholds to determine whether to actually eliminate or add candidate connections.
While, for convenience, GeNN provides some built-in models, one of GeNN’s most powerful features is that it enables users to easily define their own models for neurons, synapses and custom updates (which are typically used to implement optimizers and reductions for machine learning models [undefv]) from within Python using strings containing GeNNCode which is a subset of C99. Continuing this flexible approach, we have added support for a new model primitive – ‘custom connectivity updates’ – to GeNN which provides a framework to build structural plasticity rules. This framework is relatively complex so, while in the remainder of this section we will try to give a flavor of what is possible, we refer interested readers to the GeNN documentation (https://genn-team.github.io) for more details. Rules can be defined with the pygenn.create_custom_connectivity_update_model helper function and can access neuron and synapse variables via ‘variable references’ (pre_var_refs and post_var_refs; and var_refs respectively). Additionally, they provide one GeNNCode string which is run on the GPU and parallelized across the rows of the synaptic matrix (row_update_code) and another that is evaluated serially on the CPU (host_update_code). Within all of these code strings, a suitable random number generator is available via a family of gennrand functions. Within the row_update_code, the index of the presynaptic neuron associated with the row of connectivity can be accessed via the internal variable id_pre and synapses can be added using the add_synapse function which takes the index the postsynaptic target and the initial values of any synaptic variables, in the order they are specified in var_refs. Variables which are associated with the connectivity being updated but are not referenced in this way will be initialized to zero for the new synapse. Using this functionality, we can define a simple example which adds diagonal connections to the connectivity matrix with a synaptic weight (referenced via g) of $1.0$:
@@
At runtime, the update can be triggered (along with any other updates in the “update_connectivity” group) by simply calling model.custom_update("update_connectivity"). If row updates need to remove or otherwise access existing synapses they must iterate through them as the rows of the ragged matrix data structure are unsorted. This is implemented by the for_each_synapse statement, within the body of which, the index of current postsynaptic targets can be accessed via the id_post variable (as well as the values of any variables or variable references associated with individual synapses or postsynaptic neurons). Synapses can also be removed within the body of the for_each_synapse statement using the remove_synapse function. Finally, custom connectivity updates can define their own additional neuron and synapse variables (pre_vars and post_vars; and vars respectively). Using this additional functionality, we can define another example which, on the CPU, first loops through each presynaptic neuron and picks a random postsynaptic target to try and remove. Then, on the GPU, we search each row of connectivity in parallel and, if the chosen target is found, remove it:
@@
In order to facilitate postsynaptic spike-triggered updates to synapses, such as those involved in Spike Timing Dependent Plasticity (STDP), GeNN builds an additional transpose matrix data structure (see [undefai] for a complete description). This data structure is normally built once at initialization time but, if custom connectivity updates are applied to connections with these data structures, they get rebuilt after each custom connectivity update.
@@
While we have previously implemented complex machine learning workflows with networks of this sort in PyGeNN [undefu, undeft], the syntax is designed for computational neuroscience and it can become tedious and difficult to maintain. Therefore, we have built mlGeNN on top of GeNN to provide a Keras-like abstraction [undefaj, undefv] with support for both converting pre-trained ANNs to SNNs and for directly training SNNs using e-prop and Eventprop [undefak]. Here we have further extended mlGeNN to support DEEP R which can be invoked by creating a sparsely-connected network and ‘compiling’ it using an EPropCompiler instantiated as follows:
@@ 2.1.2 DEEP R
Figure 2: Illustration of data structures used for the DEEP R implementation. Five presynaptic neurons are connected to four postsynaptic neurons using sparse connectivity with a maximum out-degree of three. Each panel shows a network sketch, the ragged matrix and two bitfields – one mirroring the connectivity to provide a faster lookup and the other holding the fixed sign of each potential connection. (A) Initial state. Weights whose sign no longer matches the sign bit are highlighted in green. (B) Elimination pass. Connections highlighted in (A) are removed, the number of connections in each row made ‘dormant’ are counted and the bitfield is updated. (C) Formation pass. Yellow highlighting indicates new the randomly added connections.
The rewiring algorithm DEEP R [undefx] is designed for training efficient sparse neural networks by replacing weak connections with randomly sampled new ones. DEEP R operates alongside a gradient-based learning rule and marks synapses as ‘dormant’ if weight updates would have led to weights changing sign. Dormant connections are removed and an equal number of randomly chosen synapses are added as replacements. Unlike typical incremental pruning approaches, DEEP R maintains a constant level of connection sparsity throughout training. However, like many implementations of pruning in standard ML tools, the original implementation represents the sparse connectivity as an additional dense weight matrix containing the correct sign of each connection. Here we implement a more efficient version using the structural connectivity framework described in section 2.1 which only requires two additional bitfield data structures on top of the standard GeNN sparse matrix format – one to track the sign and one the presence of each potential synaptic connection.
After the standard initialization of synaptic connectivity, a custom connectivity update is used to initialize the two additional bitfields as illustrated in figure 2A. First, all connectivity bits are cleared and all sign bits are randomized to fix the signs for all potential connections. Then, the update loops through all the initialized synapses in each row, sets the corresponding connectivity bit and ensures the corresponding sign bit matches the sign of the initial weight of the actual connection. After each training batch, DEEP R is implemented by launching a sequence of custom updates.
@@
is implemented by a final custom connectivity update which first uses the CPU to sum the row-wise dormant synapse count and then re-distributes these uniformly as new ‘activations’ between the rows of synaptic connectivity. Then, on the GPU, these new synapses are created for each activation, checking the connectivity bitfield to ensure there are not duplicate synapses (figure 2B–C).
In its original form [undefx], DEEP R also features an additional noise term but, following [undefw] – who first combined DEEP R with e-prop – we do not include this in our implementation.